# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 4.1 Trade Balance
For a given period, country $i$’s trade surplus from $j$ means $j$’s trade deficit of the same magnitude from $i$, and vice versa. With a trade surplus, job opportunities are created in the country, decreasing its unemployment rate and expanding its economy of scale. It also improves the economy’s creditworthiness because it can use the surplus to pay down its foreign debt, if any. A country’s surplus could result from its increasing competitive advantage in tradable goods and services when it beats the competitors in the counterpart’s market. Opinions on the impacts of trade deficits, either bilateral or overall, are diverse and sometimes controversial, although huge deficits can undermine the sustainability of the economy. In the short run, however, trade deficits help nations avoid shortages of goods and mitigate related economic problems, such as inflation and poverty. Implications of trade deficits also depend on their impacts on national security and how they are financed. No nation would completely ignore trade deficits or competitiveness; otherwise, an one-sided emphasis will not last long.
There is a benchmark choice of $\lambda_{ji}$ for net trade balances: $\lambda_{ji}=g_{j}/g_{i}$, the ratio between their productions. In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$ while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$. If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these two countries, as claimed by Theorem 3. This condition is for instant balance change. In aggregate value over a specific time period, there is no trade surplus and deficit between these two countries if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$ because $g_{i}P_{ij}$ is the exports from $i$ to $j$ and $g_{j}P_{ji}$ the exports from $j$ to $i$. For $i$ to have a trade surplus with $j$, we need $g_{i}\lambda_{ji}\Delta P_{ji}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain this trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but hurts $\pi_{i}$ — with overly excessive exports from $i$ to $j$ in the $i$th row of $P$, country $i$ transfers additional production power to $j$, reducing $\pi_{i}$. So $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports even though $i$ loses fewer. This could eventually adversely affect the productions of both countries in the future so that most countries would avoid mutually damaging stalemates. In general, the parameter $\lambda_{ji}$ is determined by both players. However, the sign and size of $\Delta P_{ji}$ are often decided by one only, especially if it has overwhelming power over its counterpart.
@@
To ease the trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change on $\Delta P_{ji}$. On the other hand, if they look for a balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ removes any trade deficits in the future. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for a long-run rivalry due to $\pi$’s involvement in all countries and all industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and relieve relevant problems such as unemployment and national debt. According to Pozsar, (2022), for example, the 2018 China-USA trade war started in this way: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but built up massive national debt, accumulated through wars on terrors, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China wanted to build a global 5G network and make cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is a proper choice for $\lambda_{ji}$.
We study a general case when a superpower country $i$ faces imminent challenges from an emerging economy $j$. By the Matthew effect, likely $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This can happen only when $\Delta P_{ij}<0$ when both countries lose comparative advantages. Figure II illustrates this trilemma, in which up to two policy objectives are possible for country $i$. If it adopts a position on edge $b$, for example, then it upturns comparative advantage and earns a trade surplus, but gives up competitiveness unless it absorbs competitiveness from other countries. Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned with each other and also do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas $j$’s at another one. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries enjoy escalating economy of scale and diminishing marginal cost of production. According to the proof of Theorem 4, country $j$ could achieve these three objectives listed in the theorem. This seems unfair to country $i$. However, it is conditional on the Matthew effect, which is somewhat “biased” toward $\pi_{i}$. Therefore, Theorem 4 is indeed a redress to the effect.
@@
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not achieve all the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
@@
4.2 The Bargaining Solution
We consider a compromise solution between trade balance and competitiveness because a small change $\Delta P_{ji}$ would hurt or benefit both countries. In the trade system, some countries prefer competitiveness while others prefer improved trade balance (either reducing deficit or increasing surplus). Their preferences affect the choice of $\lambda_{ji}$, which holds $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If the countries are indifferent between competitiveness and trade balance, then the middle point seems to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this middle-point solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Otherwise, whenever countries $i$ and $j$ have opposite priorities between enhancing competitiveness and improving trade balance, we have a bargaining problem and apply Nash, (1950) bargaining to find a cooperative solution, which leverages between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ and also binds the condition $\lambda_{ji}\lambda_{ij}=1$.
@@ 4.2 The Bargaining Solution
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. Without $i$’s bargaining, $\lambda_{ji}$ can secure $p$ and $q$ by Corollary 1 or Theorem 3. Similarly, without $j$’s bargaining, $\lambda_{ij}$ can secure $1/p$ and $1/q$ because
$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}$

### v5

@@ 4.1 Trade Balance
For a given period, a trade surplus for country $i$ with country $j$ implies a trade deficit of the same magnitude for $j$ with $i$, and vice versa. A trade surplus generates job opportunities, reduces unemployment, and expands the economy of scale. It also enhances creditworthiness by enabling the country to pay down foreign debt. A surplus can result from competitive advantages in tradable goods and services. Opinions on trade deficits vary, though large deficits can undermine economic sustainability. In the short term, trade deficits help avoid shortages and mitigate issues like inflation and poverty. The implications of trade deficits also depend on their impact on national security and how they are financed. No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis will not endure.
There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances, $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3. For country $i$ to have a trade surplus with country $j$ due to the change $\Delta P_{ij}$, $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$. Excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$. Therefore, $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports, though $i$ loses fewer. This could adversely affect future production, so most countries avoid mutually damaging stalemates. Generally, $\lambda_{ji}$ is determined by both players, but the sign and size of $\Delta P_{ji}$ are often decided by one, especially if it has overwhelming power.
@@
To ease trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $\Delta P_{ji}$. On the other hand, if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates any future trade deficits. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for long-term rivalry due to $\pi$’s involvement in all countries and industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and alleviate related problems such as unemployment and national debt. According to Pozsar (2022), for example, the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is an appropriate choice for $\lambda_{ji}$.
We examine a scenario where a superpower country $i$ faces imminent challenges from an emerging economy $j$. According to the Matthew effect, it is likely that $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot simultaneously have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This situation occurs only when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages. Figure II illustrates this trilemma, where up to two policy objectives are achievable for country $i$. For instance, if country $i$ adopts a position on edge $b$, it can overturn comparative advantage and earn a trade surplus but sacrifices competitiveness unless it absorbs competitiveness from other countries. In the early 2000s, bilateral exports between China and the USA increased as a percentage of their respective production. According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
@@
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
@@
Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned and do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas country $j$’s interests could lie at another point. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries benefit from escalating economies of scale and diminishing marginal costs of production. According to the proof of Theorem 4, country $j$ could simultaneously achieve all three objectives listed in the theorem. This seems unfair to country $i$, but it is conditional on the Matthew effect, which is somewhat biased toward $\pi_{i}$. Therefore, Theorem 4 serves as a redress to the effect.
4.2 Nash Bargaining Solutions
This subsection examines a compromise solution between trade balance and competitiveness. In the trade system, some countries prioritize competitiveness while others aim to improve trade balance, either by reducing deficits or increasing surpluses. Their preferences influence the choice of $\lambda_{ji}$, which maintains $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If countries are indifferent between competitiveness and trade balance, the midpoint appears to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this midpoint solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Alternatively, when countries $i$ and $j$ have opposing priorities between enhancing competitiveness and improving trade balance, the Nash bargaining solution (1950) can be employed to find a cooperative solution that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while adhering to the condition $\lambda_{ji}\lambda_{ij}=1$.
@@ 4.2 Nash Bargaining Solutions
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. By Corollary 2 and Theorem 3, without bargaining, $\lambda_{ji}$ can secure $p$ and $q$, and $\lambda_{ij}$ can secure $1/p$ and $1/q$ because:
$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}\quad\mathrm{and}\quad\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
We consider two scenarios when countries $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$. By Eq. (15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other. In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$.
Theorem 5 addresses the bargaining problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is consistent across both preference scenarios, thus it remains effective when either country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. The solution remains unchanged under any positive affine transformation of their utility functions, according to a property of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation bridges the gap between these two ratios.
Theorem 5
When countries $i$ and $j$ have opposite preferences on $p$ and $q$, the Nash bargaining solution is:
$\lambda_{ji}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}\hskip 14.22636pt\mathrm{and}\hskip 14.22636pt\lambda_{ij}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}$
for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
Figure III illustrates the bargaining solution. In the left plot (first scenario), $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$ while $i$ moves from bottom to top along the same curve. When they meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, the area of the rectangle cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and the status quo point $(p,1/q)$ is maximized. The right plot (second scenario) illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences.
The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ rely on the choice of status quo points. The effective bargaining set is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$ because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second, respectively) scenario. If both players intensify or weaken their status quo positions symmetrically, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 4. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
Corollary 4
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$, the bargaining solution remains $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
Additionally, as Corollary 5 states, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$, as implied in Corollary 2 and Theorem 3. In summary, when countries $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{ll}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},&\mathrm{otherwise.}\end{array}\right.$ (16)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb.
Corollary 5
${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
Finally, country $j$ should avoid direct confrontation with a significantly larger-sized economy $i$, according to Corollary 6. This was likely China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Consequently, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, 1968).
Corollary 6
Country $j$ should avoid homogeneous preferences with a significantly larger-sized country $i$.
4.3 National Bargaining Power
In deriving the solution $\lambda_{ji}^{*}$, both countries have a symmetric role and identical prior bargaining power, as per the Nash bargaining axiom. However, countries $i$ and $j$ typically possess asymmetric prior bargaining power. For instance, we could maximize $\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}$ or $\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ given some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013). Without prior knowledge about $\zeta_{ij}$, there is no justification to favor one country over another, resulting in $\lambda^{*}_{ji}$. In a real bargaining context, however, the parameter $\zeta_{ij}$ depends on numerous characteristics between these two, as well as other, countries. Non-economic factors such as military power, population, culture, tradition, legislation, and location, along with alternative optimal criteria, can significantly compromise the symmetry axiom and lead to bargaining solutions that differ substantially from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
We introduce a posterior bargaining power that integrates both the symmetric prior and the data, capitalizing on the results in Theorem 5. To extend the bilateral Nash solution in Theorem 5 to multilateral negotiations, we define
as country $i$’s national bargaining power in the trade system. It is normalized so that all bargaining power sums to one, allowing $\beta_{i}$ to be compared over time. This measure aids multinational negotiations. For instance, when allocating votes in forming or reforming a multinational alliance, using this power can avoid bilateral dialogues and discussions. Despite factors other than production and competitiveness, bilateral trade agreements remain more prevalent than multilateral ones; Yilmazkuday and Yilmazkuday (2014) explored some reasons.
The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ constructs a linear ordering for all countries and maintains a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of $i$’s power over $j$ and $j$’s power over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$. Also, we can recover it directly from the relation $\lambda^{*}_{ik}=\beta_{i}/\beta_{k}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. Thus, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when companies in country $i$ rely on inputs from partner countries. The more companies depend on partners, the less power the country has to coerce concessions from them. From the consumption aspect, bargaining power increases when the country consumes more imported goods and services, thereby influencing the production of its counterparts. Between the determinants, $g_{i}$ represents the country’s breadth of economy, while $\pi_{i}$ predominantly reflects the depth of the economy, as the matrix $P$ in Eq. (3) already normalizes production volumes. However, the positive correlation between them implies a risk of a domino effect, where a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, leading to further reduction in $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop in $g_{i}$.
5 Economic Globalization
Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the UK’s Brexit in 2020 was a clear rejection of globalization. Conversely, protectionist measures could raise tariffs on imported goods, set import quotas for other countries, and enact stricter government regulations on imports. For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports from other countries (Washington Post, 2018).
Globalization creates winners and losers within a nation but does not automatically compensate the losers, thereby fueling income inequality and political divisions. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. The objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments in the entire matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, with a small change $\Delta P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we assume that whenever $P$ changes the $i$th column, the other columns change proportionally to ensure each row of $P$ still has a unit sum. For example, the change $\Delta P_{ji}$ causes proportional changes $\Delta P_{ji}\sigma_{ji}$ to the $j$th row where:
$\sigma_{ji}\overset{\mathrm{def}}{=\joinrel=}\frac{1}{1-P_{ji}}\left(-P_{j1},\cdots,-P_{j,i-1},1-P_{ji},-P_{j,i+1},\cdots,-P_{jn}\right).$
Consequently, the change $\Delta P_{ii}$ provokes adjustments $\Delta P$ in three levels of fallout. First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$. Secondly, the $i$th column in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}$, due to reprisal reactions, where $\odot$ denotes the element-wise product or Hadamard product and $\Lambda_{i}$ is the $i$th row of $\Lambda$. Lastly, the $i$th column of $\Delta P$ triggers changes in all other columns. We define:
$M\overset{\mathrm{def}}{=\joinrel=}\mathrm{diag}\left(\sigma_{ii}\odot\Lambda_{i}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right]$
where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with the diagonal vector $\sigma_{ii}\odot\Lambda_{i}$. Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$, and thereby $\Delta P=\Delta P_{ii}M$. So, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M$ as $\Delta P_{ii}\to 0$. We let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
Theorem 6
With the above setting and notations, we have:
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (17)